# A Survey on Adversarial Contention Resolution

### v1

@@ 1.2. Survey Organization
We use the attributes illustrated in Figure 1 to structure our review of the literature. In Section 2, we begin with contention-resolution algorithms for the static setting. This provides an introduction to the important metric of makespan, as well as the use of randomization in designing contention-resolution algorithms. Next, in Section 3, we move to dynamic model of packet arrivals, focusing on the case when arrival times are dictated by an adversary. Finally, in Section 4, we address additional adversarial behavior in the form of malicious interference on the channel.
Each section that reports on major result has the following associated components:
@@
    A table that offers a synopsis of the results in each section by organizing the results along common, pertinent attributes from Section 1.1. A brief summary of the result is also provided, which can be useful for reference purposes. Due to space constraints, these tables can be found in our electronic supplement (Section 6).
2. Static Setting
We begin our survey with the static setting for packet arrivals. This motivates a well-known class of randomized, non-adaptive backoff algorithms, with origins in decades-old networking protocols; we explore this topic in Section 2.1. Additionally, the static setting also allows us to introduce deterministic algorithms, which are newer additions to the literature; we review these results in Section 2.2.
@@ 2.1. Backoff Algorithms
Arguably, the most well-known backoff algorithm is randomized binary exponential backoff (BEB), which was introduced by Metcalf and Boggs (MetcalfeBo76, ). BEB proceeds in disjoint contention windows (CWs) consisting of consecutive slots. In each CW, a participant with a packet selects a slot uniformly at random to attempt a transmission. If successful, another packet may be processed in the same fashion, if one exists. Else, transmission is attempted in the next CW, which is double the size of the current CW. This is illustrated in Figure 2, along with pseudocode.
While originally intended for use with Ethernet over bus networks (stuck1984introduction, ), BEB has since found application in a range of domains such as wireless communication (zhou2019singletons, ; sun:backoff, ; 6859627, ; saher:log, ), transactional memory (herelihy:transactional, ; scherer:advanced, ), concurrent memory access (Ben-DavidB17, ; 80120, ; mellor-crummey:algorithms, ), and congestion control (mondal:removing, ). [^2] In practice, BEB may use a different constant-sized initial window size, and also be truncated at a maximum window size. Despite this widespread use, the makespan for BEB is sub-optimal (see Bender et al. (bender:adversarial, )); specifically, for a batch of $n$ packets that begin executing BEB at the same time, the makespan is $\Theta(n\log n)$. To understand why, we need to become familiar with the notion of contention.
@@
In order to better understand why BEB performs poorly, we consider a class of Bernoulli-backoff algorithms, where each packet sends with probability $1/w$ in each slot of a window of size $w$. Note that this differs subtly from the approaches discussed in Section 2.1, where a single slot is chosen uniformly at random in which to send. [^3] Interestingly, the makespan of Bernoulli backoff algorithms can differ significantly, as discussed in (bender:adversarial, ), but this is not critical to our discussion here. For Bernoulli-backoff algorithms it is easy to analyze the useful notion of contention (richa:jamming4, ; ChangJP19, ; BenderFGY19, ; DBLP:conf/spaa/AgrawalBFGY20, ; bender:how, ). Let $p_{i}(t)$ denote the probability that packet $i$ attempts to broadcast in slot $t$. We define the contention in slot $t$, $C(t)$, to be $C(t)=\sum_{i}p_{i}(t)$, where the sum is taken over all packets in the system at time $t$.
@@
The pseudocode for these backoff algorithms is presented in Figure 3. To the best of our knowledge, the initial CW sizes for LB, LLB, and FB are not specified exactly in the literature; $\Theta(1)$ will suffice for $LB$ and $LLB$, and $\Theta(n)$ for $FB$.
A conspicuous requirement of FB is that $n$ must be known to within a constant multiplicative factor in order to set the CW size. A number of prior works address such requirement and achieve an estimate of $n$ (DBLP:journals/jacm/GreenbergFL87, ; bender:contention, ; cali:dynamic, ; cali:design, ; bianchi:kalman, ) as a means for improving performance, by setting an initial CW size for various contention-resolution algorithms.
In terms of performance, for a batch of $n$ packets, the makespan for BEB, LB, FB, and LLB is $\Theta(n\log n)$, $\Theta\left(\frac{n\log n}{\log\log n}\right)$, $\Theta\left(n\log\log n\right)$, and $\Theta\left(\frac{n\log\log n}{\log\log\log n}\right)$, respectively (bender:adversarial, ; zhou2019singletons, ). Bender et al. also analyze polynomial backoff (PB), where the $i^{\mbox{\tiny th}}$ window has size $\Theta(i^{c})$ for some constant $c>1$. Here, the makespan is shown to be $\Theta(n/\log n)^{1+(1/c)}$.
Interestingly, all of these algorithms suffer from a throughput—the makespan divided by the number of packets $n$—that tends to zero. This observation is striking for a couple reasons. First, it illustrates the challenges raised by the static setting. In particular, the throughput for PB when $c=2$, known as quadratic backoff, is shown by Sun and Dai (sun:backoff, ) to be competitive with BEB under a saturated network. In contrast, under static arrivals, quadratic backoff has asymptotically inferior makespan.
Second, from a theoretical perspective, does a backoff algorithm exist that gives optimal throughput under the static setting? Bender et al. (bender:adversarial, ) show that, over the class of monotonic windowed backoff algorithms, LLB achieves the asymptotically-best makespan  (bender:adversarial, ). In other words, a non-monotonic approach is required to achieve $\Theta(n)$ makespan or, equivalently, $\Theta(1)$ throughput. While we will later revisit the first two points, it is this question of optimality that we pursue next.
@@
Ultimately, for the static setting, the above algorithms fall short due to their inability to provide $\Theta(1)$ contention over $\Omega(n)$ slots. Next, we examine algorithms that do satisfy this property.
@@
Based on our observations in Section 2, any algorithm that hopes to achieve $O(n)$ makespan must incorporate a mechanism for maintaining constant contention for $\Omega(n)$ slots. Sawtooth-Backoff (STB) (GreenbergL85, ; Gereb-GrausT92, ) accomplishes this; it is non-monotonic algorithm that executes over a doubly-nested loop. The outer loop sets the current window size $w$ to be double the one used in the preceding outer loop; this is like BEB. Additionally, for each such window, the inner loop executes over $\lg w$ windows of decreasing size: $w,w/2,w/4,...,1$. For each such window, a slot is chosen uniformly at random for the packet to transmit; this is the “backon” component of STB. When the outer loop achieves a window of size $cn$ for a sufficiently large constant $c>0$, with high probability all packets will succeed in the corresponding inner loop. Intuitively, this is because the backon component maintains constant contention in each slot of the inner-loop windows (until all packets succeed). For a single batch of $n$ packets, the makespan of STB is $O(n)$ (Gereb-GrausT92, ; GreenbergL85, ; bender:adversarial, ; zhou2019singletons, ).
A variant called Truncated Sawtooth-Backoff (TSTB) (bender:contention, ) executes identically to STB except that, for some constant $c>0$, the inner loop executes over $\lg(c\lg w)$ windows of decreasing size: $w,w/2,w/4,...,\max\{\lfloor w/c\lg n\rfloor,1\}$. Another asymptotically-optimal result is given by Anta et al. (FernandezAntaMoMu13, ; fernandez-anta:unbounded, ) with their algorithm One-fail Adaptive, which succeeds with probability $1-O(1/n)$, again for the case with no-CD and no knowledge of system size. The authors tightly analyze both One-fail Adaptive and STB, yield respective makespan guarantees of $2(\delta+1)n+O(\log^{2}n)$ and $4(1+1/\delta)n$, where $\delta$ is a small positive constant.
Anta and Mosteiro (Anta2010, ) look at a static case—referred to as static $k$-selection (see Section 3.2)—where an unknown $k$ out of $n$ participants must each succeed. For an error parameter $\epsilon\leq 1/(n+1)$, the authors give a randomized algorithm that with probability $1-O(\epsilon)$ has makespan $O(k+\log^{2}(1/\epsilon))$; for consistency, we report the case for $k=n$ in Table 1 (see our Electronic Supplements in Section LABEL:xyz).
Wang (dingyu:optimal, ) provides exact, optimal results for the special case of $n=2$. Specifically, the authors examine the minimum: average latency until both participants succeed, the latency until the first participant succeeds, and the latency until the second (last) device succeeds; the expected values for these metrics are shown to be $\sqrt{3/2}+3/2$, $2$ and approximately $3.33641$.
These results settle the question of optimal makespan for the static setting. However, an obvious question is whether this optimality holds when packets may arrive over time, and we explore this issue in Section 3.
@@ 2.2. Deterministic and Hybrid Approaches
Discussion. For the static case, many randomized algorithms have been proposed, with STB providing one of the strongest with-high-probability results: asymptotically-optimal makespan with no-CD, no knowledge of $n$, and only $O(\log n)$ channel-access attempts.
@@
Finally, in the context of network traffic, we highlight the underlying assumption that each transmission of a packet is accommodated by a single slot. Unsurprisingly, BEB also has sub-optimal makespan for the more general case with heterogeneous packet sizes, where a packet of size $k$ must have sole access to the channel for $k$ consecutive slots in order to succeed (bender:heterogeneous, ). Along similar lines, in the wireless domain, it has been argued that packet size should be incorporated into the definition of makespan, since collisions tend to consume time proportional to packet size (anderton:is, ; anderton:windowed, ).
@@ 3. Dynamic Arrivals
However, modeling of arrival times via a stochastic process may not always be justified. In particular, bursty traffic is common in practice (williamson:internet, ; wilson1991high, ; paxson:wide-area, ; Teymori2005, ; yu:study, ), and its impact has been examined (Ghani:2010, ; sarkar:effect, ; canberk:self, ; bhandari:performance, ). Indeed, the static setting discussed in Section 2 illustrates the impact on makespan by bursts of packets, where the packets of each burst succeed before the next burst arrives. This has led the research community to consider more general arrival models.
Adversarial queuing theory (AQT) (borodin2001adversarial, ; cruz-one, ; cruz-two, ; andrews2001universal, ) has been used to model packet arrivals times, and many contention-resolution results fit within the AQT framework; that is, they address adversarial models that are parameterized by rates of packet injection and burstiness. Here, a key metric is stability, which we discuss in Section 3.1.

### v3

@@ 1.2. Survey Organization
We use the attributes discussed above and illustrated in Figure 1 to structure our review of the literature. In Section 2, we begin with contention-resolution algorithms for the static case. This provides an introduction to the important metric of makespan, as well as the use of randomization in designing contention-resolution algorithms. Next, in Section 3, we move to dynamic model of packet arrivals, focusing on the case when arrival times are dictated by an adversary. Finally, in Section 4, we address additional adversarial behavior in the form of malicious interference on the channel.
Each section has the following associated components:
@@
    A table that offers a synopsis of the results in each section by organizing the results along common, pertinent attributes from Section 1.1. A brief summary of the result is also provided, which can be useful for reference purposes. Due to space constraints, these tables can be found in our electronic supplement.
2. The Static Case
We begin our survey with the static case for packet arrivals. This motivates a well-known class of randomized, non-adaptive backoff algorithms, with origins in decades-old networking protocols; we explore this topic in Section 2.1. Additionally, the static case also allows us to introduce deterministic algorithms, which are newer additions to the literature; we review these results in Section 2.2. Finally, we discuss the closely-related wakeup problem in Section 2.3.
@@ 2.1. Backoff Algorithms
Arguably, the most well-known backoff algorithm is binary exponential backoff (BEB), which was introduced by Metcalf and Boggs (MetcalfeBo76, ), and briefly discussed in Section 1. BEB proceeds in disjoint contention windows (CWs) consisting of consecutive slots. In each CW, a participant with a packet selects a slot uniformly at random to attempt a transmission. If successful, another packet may be processed in the same fashion, if one exists. Else, transmission is attempted in the next CW, which is double the size of the current CW. This is illustrated in Figure 2, along with pseudocode.
While originally intended for use with Ethernet over bus networks (stuck1984introduction, ), BEB has since found application in a range of domains such as wireless communication (zhou2019singletons, ; sun:backoff, ; 6859627, ; saher:log, ), transactional memory (herelihy:transactional, ; scherer:advanced, ), concurrent memory access (Ben-DavidB17, ; 80120, ; mellor-crummey:algorithms, ), and congestion control (mondal:removing, ). [^3] A ubiquitous example is WiFi, where many of the standards use BEB as a core mechanism for governing how devices gain access to a wireless channel. In these standards, the initial window size is typically some constant larger than $1$, and the maximum window size is limited to a value just over $1000$. Despite this widespread use, the makespan for BEB is sub-optimal (see Bender et al. (bender:adversarial, )); specifically, for a batch of $n$ packets that begin executing BEB at the same time, the makespan is $\Theta(n\log n)$. To understand why, we need to become familiar with the notion of contention.
@@
In order to better understand why BEB performs poorly, we consider a class of Bernoulli-backoff algorithms, where each packet sends with probability $1/w$ in each slot of a window of size $w$. Note that this differs subtly from the approaches discussed in Section 2.1, where a single slot is chosen uniformly at random in which to send. For Bernoulli-backoff algorithms it is easy to analyze the useful notion of contention (richa:jamming4, ; ChangJP19, ; BenderFGY19, ; DBLP:conf/spaa/AgrawalBFGY20, ; bender:how, ). Let $p_{i}(t)$ denote the probability that packet $i$ attempts to broadcast in slot $t$. We define the contention in slot $t$ to be $C(t)$ $=\sum_{i}p_{i}(t)$; that is, the sum of the sending probabilities for all packets in the system at time $t$.
@@
The pseudocode for these backoff algorithms is presented in Figure 3. To the best of our knowledge, the initial CW sizes for LB, LLB, and FB are not specified exactly in the literature; $\Theta(1)$ will suffice for $LB$ and $LLB$. A conspicuous requirement of FB is that $n$ must be known to within a constant multiplicative factor in order to set the CW size. A number of prior works show how to get an estimate of $n$ (DBLP:journals/jacm/GreenbergFL87, ; bender:contention, ; cali:dynamic, ; cali:design, ; bianchi:kalman, ).
In terms of performance, for a batch of $n$ packets, the makespan for BEB, LB, FB, and LLB is $\Theta(n\log n)$, $\Theta\left(\frac{n\log n}{\log\log n}\right)$, $\Theta\left(n\log\log n\right)$, and $\Theta\left(\frac{n\log\log n}{\log\log\log n}\right)$, respectively (bender:adversarial, ; zhou2019singletons, ). Bender et al. also analyze polynomial backoff (PB), where the $i^{\mbox{\tiny th}}$ window has size $\Theta(i^{c})$ for some constant $c>1$. The window scaling function is slightly different here; ignoring constants, we solve for $f(i)$ in $(i+1)^{c}=(1+f(i))i^{c}$, implies that $f(i)=((1+i)^{c}/i^{c})-1=((i+1)/i)^{c}-1$. For PB, Bender et al. show that the makespan is $\Theta(n/\log n)^{1+(1/c)}$.
Interestingly, all of these algorithms suffer from a throughput that tends to zero. This observation is striking for a couple reasons. First, it illustrates the challenges raised by the static case. In particular, the throughput for PB when $c=2$, known as quadratic backoff, is shown by Sun and Dai (sun:backoff, ) to be competitive with BEB under a saturated network. In contrast, under static arrivals, quadratic backoff has asymptotically inferior makespan.
Second, from a theoretical perspective, does a backoff algorithm exist that gives optimal throughput under the static case? Bender et al. (bender:adversarial, ) show that, over the class of monotonic windowed backoff algorithms, LLB achieves the asymptotically-best makespan  (bender:adversarial, ). In other words, a non-monotonic approach is required to achieve $\Theta(n)$ makespan or, equivalently, $\Theta(1)$ throughput. This question of asymptotic optimality is what we pursue next.
@@
Ultimately, for the static case, the above algorithms fall short due to their inability to provide constant contention over $cn$ slots for some sufficiently large constant $c$. Next, we examine algorithms that do satisfy this property.
@@
Based on our observations in Section 2, any algorithm that hopes to achieve $O(n)$ makespan must incorporate a mechanism for providing constant contention over $\Omega(n)$ slots. Sawtooth-Backoff (STB) (GreenbergL85, ; Gereb-GrausT92, ) accomplishes this; it is non-monotonic algorithm that executes over a doubly-nested loop. The outer loop sets the current window size $w$ to be double the one used in the preceding outer loop; this is like BEB. Additionally, for each such window, the inner loop executes over $\lg w$ windows of decreasing size: $w,w/2,w/4,...,1$. For each such window, a slot is chosen uniformly at random for the packet to transmit; this is the “backon” component of STB. When the outer loop achieves a window of size $cn$ for a sufficiently large constant $c>0$, with high probability all packets will succeed in the corresponding inner loop. Intuitively, this is because the backon component maintains constant contention in each slot of the inner-loop windows (until all packets succeed). For a single batch of $n$ packets, the makespan of STB is $O(n)$ (Gereb-GrausT92, ; GreenbergL85, ; bender:adversarial, ; zhou2019singletons, ).
A variant called Truncated Sawtooth-Backoff (TSTB) (bender:contention, ) executes identically to STB except that, for some constant $c>0$, the inner loop executes over $\lg(c\lg w)$ windows of decreasing size: $w,w/2,w/4,...,\max\{\lfloor w/c\lg n\rfloor,1\}$. Another asymptotically-optimal result is given by Anta et al. (FernandezAntaMoMu13, ; fernandez-anta:unbounded, ) with their algorithm One-fail Adaptive, which succeeds with probability $1-O(1/n)$, again for the case with no-CD and no knowledge of system size. The authors tightly analyze both One-fail Adaptive and STB, yielding respective makespan guarantees of $2(\delta+1)n+O(\log^{2}n)$ and $4(1+1/\delta)n$, where $\delta$ is a small positive constant.
Anta and Mosteiro (Anta2010, ) look at a formulation of the static case—referred to as static $k$-selection (see Section 3.2)—where an unknown $k$ out of $n$ participants must each succeed. For an error parameter $\epsilon\leq 1/(n+1)$, the authors give a randomized algorithm that with probability $1-O(\epsilon)$ has makespan $O(k+\log^{2}(1/\epsilon))$; for consistency, we report the case for $k=n$ in Table 1 of our electronic supplement.
For the special case where $n=2$, Wang (dingyu:optimal, ) provides exact, optimal results. Specifically, the average latency until both participants succeed, the latency until the first participant succeeds, and the latency until the second (last) device succeeds are shown to have expected values, respectively, of $\sqrt{3/2}+3/2$, $2$ and approximately $3.33641$.
These results settle the question of optimal makespan for the static case. However, an obvious question is whether this optimality holds when packets may arrive over time, and we explore this issue in Section 3.
@@ 2.2. Deterministic and Hybrid Approaches
2.3. The Wakeup Problem.
We introduce the wakeup problem, which is described as follows. There are $n$ participants, of which some are active (or “awake”), and the remainder are inactive (or“sleeping”) with only the ability to listen. Each awake participant has a packet to send, and upon the first success, all sleeping participants “wake up” and the problem is solved. Typically, the metric of interest is the number of rounds until the first success. Thus, contention resolution differs from wakeup: it requires all participants to succeed, while wakeup hinges on just the first success.
Despite this difference, historically the wakeup problem has considerable overlap in the literature with contention resolution, and is sometimes referred to by the same name (e.g., (fineman:contention, ; fineman:contention2, ; gilbert2021contention, ; DEMARCO20171, ; de2013contention, )). To continue the naming complications, in the static case where the set of active participants are fixed from the start, the wakeup problem is sometimes referred to as leader election (LE).
Static wakeup and LE may be viewed as a natural starting point for solving contention resolution, as repeatedly applying them over all $n$ packets offers a solution, albeit an inefficient one. A more meaningful connection can be drawn between static wakeup and contention resolution: wake-up protocols often produce an estimate for the $n$, the size of the system, which may be used to bootstrap a contention resolution protocol. More specifically, many algorithms are “uniform” or ‘fair”: in any fixed round, each participant sends packets with the same probability (independently of the prior communication history). For such algorithms, the reciprocal of the sending probability when the first success occurs may yield an accurate estimate of $n$; obtaining system-size knowledge is a well-known component of contention resolution (e.g., (jurdzinski:energy, ; chang:exponential-jacm, ; gilbert2021contention, )). For instance, such an estimate can be used to set the first window size in algorithms such as FB and STB/TSTB (recall Section 2.1), bypassing much of the backing off that would otherwise be required, and speeding up the runtime by a constant factor.
For static wakeup, makespan refers to number of slots until the first success (recall Section 1.1.3). An early result on the wakeup problem is by Willard (willard:loglog, ), who considers the case with CD. Here, Willard establishes a lower bound on uniform algorithms—for any fixed round, each participant sends packets with the same probability (independently of the prior communication history)—showing that the expected makespan is at least $\lfloor\lg\lg n\rfloor-O(1)$, and an upper bound of $\lg\lg n+O(1)$ for expected makespan is also given.
In the case of no-CD and where $n$ is known, Kushilevitz and Mansour (kushilevitz1998omega, ; kushilevitz1993omega, ) prove a lower bound of $\Omega(\log n)$ expected makespan; this applies even to unfair algorithms. An asymptotically-matching upper bound on expected makespan, and a bound of $O(\log^{2}n)$ w.h.p. is achieved by Bar-Yehuda et al. (bar1992time, ). In subsequent work that unifies many prior lower-bound arguments, Newport (newport2014radio, ; newport:radio-journal, ) also shows $\Omega(\log n)$ expected makespan and $\Omega(\log^{2}n)$ makespan w.h.p. for any randomized algorithms; these bounds hold even with multiple channels and/or CD. For algorithms that succeed with probability at least $1-\epsilon$, this lower bound was strengthened to $\Omega(\log n\,\log(1/\epsilon))$ by Farach-Colton et al. (farach-colton:lower-bounds, ). We note a couple things. First, these lower bounds are dramatically different from the lower bound by Willard (willard:loglog, ), where CD is available. Second, Jurdziński and Stachowiak (jurdzinski2002probabilistic, ) provide an algorithm with $O(\log n\,\log(1/\epsilon))$ makespan so long as at least one of the following is true: devices have unique labels, or $n$ is known.
In a setting with $\mathcal{C}>1$ channels and CD, Fineman et al. (fineman:contention2, ) give an algorithm guaranteeing w.h.p. $O((\log n/\log\mathcal{C})$ $+(\log\log n)(\log\log\log n))$ makespan for the case where any number of participants may start in the active state, along with an algorithm guaranteeing w.h.p. $O((\log n/\log\mathcal{C})+(\log\log n))$ makespan in the restricted case where at most two participants become active; this second case matches the lower bound provided by Newport (newport2014radio, ; newport:radio-journal, ). The authors also note that, while their result is described for the static case, it holds for the dynamic case.
The signal-to-interference-plus-noise ratio (SINR) model has been considered in the context of several fundamental wireless-communication problems (for details, see (moscibroda2006complexity, ; jurdzinski2012distributed, ; jurdzinski2013distributed, ; moses2020deterministic, ; jurdzinski2015cost, ; JurdzinskiR17, )). Informally, under the SINR model, the ratio of (i) the power of an incoming transmission to (ii) the sum interference from other devices plus the floor noise, must exceed a threshold value in order for the transmission to succeed. In this setting, Fineman et al. (fineman:contention, ) shows how to solve wakeup with $O(\log n+\log R)$ makespan w.h.p., where $R$ is ratio between the longest and shortest link. The authors point out that, in many practical settings, $R$ is at most a polynomial in $n$, which yields an $O(\log n)$ result. [^5] This ratio appears in other related results under the SINR model, such as interference scheduling or link scheduling (e.g., (fanghanel2011improved, )). The authors also show a general lower bound (e.g., it is not restricted to the class of fair algorithms) of $\Omega(\log n)$.
Finally, we note a recent result by Gilbert et al. (gilbert2021contention, ), that examines the implications of system-size “predictions”; that is, the algorithm is given access to the probability distribution defined over the possible values of $n$. The authors give lower and upper bounds for both deterministic and randomized solutions, for the CD and no-CD cases. Interestingly, in addition to new bounds, this work recovers some lower bounds, including the $\Omega(\log\log n)$ lower bound by Willard (willard:loglog, ) mentioned above.
@@
Summary. For static contention resolution, many randomized algorithms have been proposed, with STB providing one of the strongest with-high-probability results: asymptotically-optimal makespan with no-CD, no knowledge of $n$, and only $O(\log n)$ channel-access attempts.
@@
In the context of network traffic, we highlight the underlying assumption that each transmission of a packet is accommodated by a single slot. Unsurprisingly, BEB also has sub-optimal makespan for the more general case with heterogeneous packet sizes, where a packet of size $s$ must have sole access to the channel for $s$ consecutive slots in order to succeed (bender:heterogeneous, ). Along similar lines, in the wireless domain, it has been argued that packet size should be incorporated into the definition of makespan, since in WiFi networks collisions may consume time proportional to packet size (anderton:is, ; anderton:windowed, ).
For static wakeup, under each of CD or no-CD, asymptotically matching upper and lower bounds are known. It is interesting to compare these results to those of static contention resolution. In particular, while the makespan for contention resolution is higher, the optimal average per-packet time to success is $\Theta(1)$, which is lower than even the CD case for wakeup. As we will see, this is also a common theme in the dynamic case with randomized approaches.
@@ 3. Dynamic Arrivals
However, modeling of arrival times via a stochastic process may not always be justified. In particular, bursty traffic is common in practice (williamson:internet, ; wilson1991high, ; paxson:wide-area, ; Teymori2005, ; yu:study, ), and its impact has been examined (Ghani:2010, ; sarkar:effect, ; canberk:self, ; bhandari:performance, ). Indeed, the static case discussed previously in Section 2 illustrates the impact on makespan by disjoint bursts of packets. This has led the research community to consider more general arrival models.
To this end, adversarial queuing theory (AQT) (borodin2001adversarial, ; cruz-one, ; cruz-two, ; andrews2001universal, ) has been used to model packet arrivals times, and many contention-resolution results fit within the AQT framework; that is, they address adversarial models that are parameterized by rates of packet injection and burstiness. Here, a key metric is stability, which we discuss in Section 3.1.